OpenStudy (anonymous):

How to solve this differential equation: x*2 dy/dx = e*y

OpenStudy (unklerhaukus):

separate the variables multiply by dx, divide by x^2, divide by e^y....

OpenStudy (anonymous):

How do I separate the variables ?

OpenStudy (unklerhaukus):

follow those steps

OpenStudy (unklerhaukus):

\[x^2 \frac{\text dy}{\text dx} = e^y\] multiplying by \(\text dx\) \[x^2 \text dy = e^y\text dx\] dividing by \(x^2\) \[ \text dy = e^yx^{-2}\text dx\] dividing by \(e^y\) \[ e^{-y}\text dy =x^{-2}\text dx\] now you can integrate \[ \int e^{-y}\text dy =\int x^{-2}\text dx\]

OpenStudy (anonymous):

Is the answer ln x +c ?

OpenStudy (unklerhaukus):

\[-e^{-y}=-x^{-1}+c_1\]\[e^{-y}=x^{-1}+c\]\[\ln{e^{-y}}=\ln(x^{-1}+c)\]\[-y=\ln{(x^{-1}+c)}\]\[y=-\ln (x^{-1}+c)\]

OpenStudy (anonymous):

I missed the minus sign. Thanks.

OpenStudy (unklerhaukus):

yeah you were close